OpenStudy (anonymous):

Will give medal! April adds the expressions 9x + 7 and 5x – 3. Evan adds the expressions 14x + 2 and 4x + 3. How much greater is Evan’s sum than April’s sum?

OpenStudy (anonymous):

Well lets add April's terms 9x+7+5x-3

OpenStudy (anonymous):

collect like terms... 9x+5x=? 7-3?

OpenStudy (anonymous):

7-3=4 and 9x+5x=?

OpenStudy (anonymous):

so april's sum is 14x+4 now lets look at evan's

OpenStudy (anonymous):

14x+2+4x+3=?

OpenStudy (anonymous):

14x and 4x are tike terms and we have 14x+4x which equals? and 2+3=?

OpenStudy (im_an_angel):

the first equation when added is 14x+4 and the second is 18x+5 so subtract hem now :)

OpenStudy (anonymous):

Thx